OpenStudy (anonymous):

can anyone help me with better understanding, How to solve linear equations by graphing? thanks!

Mehek (mehek14):

do you have an attachment?

OpenStudy (anonymous):

yea you need an attachment

Mehek (mehek14):

the formula for linear equations is: \(\color{blue}{y = mx + b}\)

OpenStudy (jdoe0001):

http://www.mathwarehouse.com/algebra/linear_equation/slope-of-a-line.php <--- find the slope, and that slope will be "ratio" that y is to x and where the lines "touches" or "intercepts" y, is the well, y-intercept thus y = [slope] x + [y-intercept]

OpenStudy (anonymous):

Attachment meaning?

OpenStudy (anonymous):

picture to go with it

OpenStudy (anonymous):

so how would I solve this linear equation X+3y=8 x-4y=-6 by graping,

Mehek (mehek14):

turn it into y = mx + b

OpenStudy (anonymous):

so it would be (-x+8)

OpenStudy (anonymous):

y=mx+b for this equation would be (-y)=(-x+8) right?

OpenStudy (bee_see):

for x+3y=8? no, it's not -y

OpenStudy (anonymous):

How do I find y=mx+b?

OpenStudy (bee_see):

you want to move x to the right side. when you move it, the sign of it will change.

OpenStudy (anonymous):

so it would change to y?

OpenStudy (bee_see):